Please help me solve this question
The pressure P of a certain gas is related to volume V according to: P=100/V
(a) Is 0 in the domain of this function?
(b) What are P(100) and P(50)?
(c) As volume decreases, what happens to pressure?

Answer by Alan3354(69443)

The pressure P of a certain gas is related to volume V according to: P=100/V
P(V) = 100/V
---
(a) Is 0 in the domain of this function?
No.
(b) What are P(100) and P(50)?
P(100) = 100/100 = 1
----
Do P(50) the same way.
---
(c) As volume decreases, what happens to pressure?
Look at some values and see.